### v1

@@ Abstract
The Giant Radio Array for Neutrino Detection (GRAND) is a proposed multi-messenger observatory of ultra-high-energy (UHE) particles of cosmic origin. Its main goal is to find the long-sought origin of UHE cosmic rays by detecting large numbers of them and the secondary particles created by their interaction—gamma rays, and, especially, neutrinos. GRAND will do so using large arrays of radio antennas that look for the radio signals emitted by the air showers initiated by the interactions of the UHE particles in the atmosphere. Since 2023, three small-scale prototype GRAND arrays have been in operation: GRAND@Nançay in France, GRAND@Auger in Argentina, and GRANDProto300 in China. Together, their goal is to validate the detection principle of GRAND under prolonged field conditions, achieving efficient, autonomous radio-detection of air showers. We describe the hardware, software, layout, and operation of the GRAND prototypes and show the first radio spectra measured by them. Despite challenges, the successful operation of the prototypes confirms that the GRAND instrumentation is apt to address the goals of the experiment and lays the groundwork for its ensuing stages.
@@ 1 Introduction
The Giant Radio Array for Neutrino Detection (GRAND) [1] is a planned multi-messenger observatory of ultra-high energy (UHE) particles of cosmic origin—cosmic rays, neutrinos, and gamma rays. Its ultimate goal is to provide crucial evidence that addresses the long-standing question of the origin of UHE cosmic rays (UHECRs) [2, 3, 4]. GRAND plans to do so, directly, by detecting a vast number of UHECRs and, indirectly, by detecting the secondary gamma rays and, especially, the UHE neutrinos made by the interaction of UHECRs [5]. To achieve this, GRAND plans to use large arrays of radio antennas that search for the distinctive radio signals emitted by the extensive air showers initiated by UHE particles interacting in the atmosphere [6, 7].
@@
The radio-detection of UHECRs is a mature technique. It has been demonstrated in past and present detectors using overground antenna arrays [6, 7] by the Auger Engineering Radio Array (AERA) [8], the Low-Frequency Array (LOFAR) [9], the Taiwan Astroparticle Radiowave Observatory for Geo-synchrotron Emissions (TAROGE) [10], and the Tianshan Radio Experiment for Neutrino Detection (TREND) [11]; using underground arrays [12] by the Askaryan Radio Array (ARA) [13] and the Antarctic Ross Ice Shelf Antenna Neutrino Array (ARIANNA) [14]; and using balloon flights by the Antarctic Impulsive Transient Antenna (ANITA) [15]. Concurrent with GRAND, the new Radio Detector of the AugerPrime upgrade [16], the upcoming Payload for Ultrahigh Energy Observations (PUEO) [17], and the proposed next-generation Beamforming Elevated Array for COsmic Neutrinos (BEACON) [18] and Radar Echo Telescope (RET-CR) [19] also target the radio-detection of UHE particles, though adopting different, complementary [20] detection strategies. References [21, 22, 23] present an overview of upcoming UHE detectors.
@@
Figure 1: The GRAND prototype arrays. The three GRAND prototype arrays were deployed in 2023 and have been in operation since. They share a common basic instrumentation design, modified to suit the local environmental condition at the array site, and to test the performance of alternative hardware and software choices. See the main text for details. (Credit for the base map: Wikimedia Commons, public domain.)
@@
Building on the experience gained by these experiments—most directly, by AERA and TREND—GRAND plans to deploy antenna arrays at a larger scale [1, 24]. This involves increasing the number of antennas deployed—to eventually tens of thousands—and deploying them in arrays that are sparser than before—with km-scale antenna separations. This up-scaling is requisite for the antenna arrays to monitor the large volumes of the atmosphere needed to deliver vast numbers of detected UHECRs [1] and a sensitivity to even tiny fluxes of UHE neutrinos [1, 25, 22, 21, 23]. The GRAND Collaboration plans to achieve this while keeping its carbon footprint low [26].
Such an enterprise involves facing technical challenges that, while significant, are surmountable. To face them, GRAND is to be built with a staged construction plan [1, 27, 28] designed to validate the design and operation of the experiment in smaller scales before deploying larger arrays.
@@
Figure 1 shows the locations of the GRAND prototypes, in operation since 2023 [28]: GRAND@Nançay in Nançay, France, GRANDProto300 [29] in Dunhuang, China, and GRAND@Auger [30] in Malargüe, Argentina, at the site of the Pierre Auger Observatory (Pierre Auger) [31]. Their purpose is to validate two core requirements that GRAND must fulfill to achieve its goals.
@@
First, the prototypes must validate the viability of autonomous radio-detection of air showers in a sparse radio array [32, 33, 34, 35], unaided by particle detectors, and despite the presence of natural and anthropogenic radio background that may mimic the radio signatures expected from showers. Second, the prototypes must validate that the arrays are able to operate under harsh environmental field conditions—high temperature, sandstorms, rainstorms—for prolonged periods of time. The lessons garnered from the prototypes will inform the design of the experiment in later stages.
In this paper, we introduce the design and operation of the GRAND prototype arrays, focusing on the two larger ones, GRANDProto300 (GP300) and GRAND@Auger (G@A). In addition, we present the first results obtained by them during their first two years of operation: the time-dependent and time-averaged radio spectra measured by them, including identified sources of radio background. The ongoing operation of the arrays—though not devoid of challenges—coupled with their early-data results, validate the resilience of the GRAND design to adapt to different environmental conditions, a requirement to pursue future, larger stages of GRAND. This constitutes the basis for demonstrating the autonomous detection of cosmic rays in GRAND, which will be shown in future publications (see Ref. [35] for preliminary results).
@@
The remainder of the paper is organized as follows. Section 2 introduces the GRAND prototype arrays, their sites, and layouts. Section 3 describes the GRAND detection unit, its hardware, software, and performance. Section 4 presents the self-trigger algorithm of the experiment and the data acquisition software. Section 5 describes the data transfer pipeline, from local and central storage on-site to permanent storage off-site. Section 6 presents the offline data-quality monitoring tools of the experiment. Section 7 shows the first radio spectra observed by the GRAND prototypes, including identified background sources. Section 8 summarizes and concludes.
@@ 2 Overview of the GRAND prototype arrays
Figure 2: Array layout of GRANDProto300 (left) and of GRAND@Auger (right). Each GRAND detection unit (DU) has an assigned identification number. The Central Station in GRANDProto300 and the Central Radio Station in GRAND@Auger house the central data acquisition (DAQ) systems of the experiments. For GRANDProto300, the DUs from the 13-antenna stage, GP300_s13, were repositioned as part of the 65-antenna stage, GP300_s65, in May of 2025, and the original Central Station 1 (CS1) was replaced by the new CS2. Positions are given with respect to DU 1 of GP300_s65. For GRAND@Auger, we show also the positions of nearby AERA radio-stations and access roads. Positions are given in the Auger Coordinate System. See Section 2 for details.
The site and array layout for each GRAND prototype were decided to fulfill their specific goals: testbenching for GRAND@Nançay, demonstrating the autonomous radio-detection of air showers for GP300, and observing coincident detection of air showers with the Pierre Auger detectors in G@A. The radio background varies between sites, with the background at GRAND@Nançay being the loudest, and the background at GP300 being appreciably smaller than at G@A, predominantly due to GP300 being farther from human activity.
GRAND@Nançay.—GRAND@Nançay is a small array of four GRAND detection units (DUs, Section 3) located at the Nançay Radio Observatory, in France. It was conceived as a field site accessible to European laboratories, facilitating testbenching of equipment and new ideas in design, trigger, and data acquisition. The design of GRAND@Nançay resembles that of the other two prototype arrays, though with notable differences, such as having antennas of lower height and communications via optical fiber instead of wireless. Due to its small size, it serves primarily as a testbed. In the remainder of the paper, we focus on GRANDProto300 and GRAND@Auger.
GRANDProto300.—GP300 is the pathfinder array for GRAND in the Northern Hemisphere. In its final stage, it will be a 300-antenna array; presently, it exists in partial configurations.
@@
In June of 2023, the deployment of the first, 13-antenna stage, GP300_s13, was completed [36]. During its nearly two years of operation, it enabled the validation of the GRAND detection concept and the improvement of the system design. By April of 2025, the array was extended to its current 65-antenna stage, GP300_s65, which has been operating since and whose additional goals include testing the efficiency and purity of the detection of inclined air showers from UHECRs. GP300, even in partial configuration, is expected to detect UHECRs with energies between 100 PeV and a few EeV, probing the Galactic to extragalactic transition region of cosmic ray origin.
@@
Figure 2 shows the layout of GP300_s13 and GP300_s65. In GP300—and also in G@A—DUs are arranged in a hexagonal layout, a choice made to optimize the simulated efficiency of air-shower detection with respect to the density of antennas on the ground. A hexagonal array has fewer directions along which the performance of shower-reconstruction methods is degraded as a result of the shower trajectory being aligned with the axes of symmetry of the array layout.
@@
GP300_s13 had a three-hexagon layout, covering an area of about 2 km2 with a DU spacing of about 500 m. The Central Station 1 (CS1) was built north-west of the array; it consisted of three sets of container rooms that stored the central data acquisition (DAQ) computer and the lodging and cooking facilities. GP300_s13 was decommissioned in April of 2025 and its DUs were incorporated into GP300_s65, which has remained operational since. GP300_s65 was built south of GP13, covering an area of about 30 km2. The array has a dense region to the east with a DU spacing of 577 m and a sparse region to the west with a spacing of 1000 m. A new Central Station (CS2) was built for GP300_s65 and placed at its center.
@@
Preliminary simulations [37] estimate that the full GP300 array may detect about 130 inclined cosmic-ray events per day—for an ideal 100% uptime of the detector—in the energy range of $10^{17}$$10^{20}$ eV, reaching a trigger efficiency of 90% already at $10^{18.3}$ eV. These estimates account for the processing of the radio signal from the air shower through the radio-frequency chain of the detector units (more on this later) and for the persistent background noise from the Galactic radio emission. Improvement are underway to optimize the cosmic-ray trigger rate, suppress radio background, and refine the selection of cosmic-ray candidates [38, 39, 40, 41, 42, 43]. Because GP300_s13 and GP300_s65 are smaller, they are sensitive to the lower end of the cosmic-ray energy range, between about $10^{16.5}$ eV and $10^{18.5}$ eV, where the cosmic-ray flux is higher. Although the search for these cosmic rays is complicated by firmware and DAQ changes made during prototype testing, it has already yielded its first likely candidates in GP300_s65 data [38, 39, 40].
@@
GRAND@Auger.—G@A results from an agreement between the GRAND and Pierre Auger Collaborations to repurpose 10 AERA radio-stations in Malargüe, Argentina, into GRAND DUs. The deployment of G@A was completed in August of 2023 and it has been operating since. Because G@A is embedded within the region of the Pierre Auger detector array, it offers, uniquely among the three prototypes, the opportunity to validate the radio-detection—and ensuing reconstruction—of air showers that are simultaneously detected by G@A and by the Pierre Auger surface detector (SD) array of water-Cherenkov particle detectors, fluorescence telescopes, and radio antennas, the latter recently deployed [16].
@@
Figure 2, right, shows the layout of G@A, deployed within the AERA array, itself located within the Pierre Auger 750-m SD array, and within the field of view of the Coihueco and HEAT fluorescence detectors [44, 45]. Ten of the AERA Phase-2 stations were modified into GRAND DUs. The DUs are laid on two superimposed hexagons with a predominant spacing of 250 m, covering an area of 0.5 km2. The G@A central DAQ computer is located in the AERA Central Radio Station (CRS). Nearby roads facilitate access to the site.
@@
For G@A, using a denser detector layout enhances the detection of less energetic and less inclined events. The estimated event rate for the Pierre Auger 750-m SD array is $390\pm 17$ showers per day over 25 km2 [46]. This corresponds to about 8 showers per day detected by the Pierre Auger SD array within the area of G@A, making it viable to search for showers measured simultaneously by G@A and Pierre Auger detector arrays. One such coincident event has been detected so far [47].
Because the geomagnetic field is weaker at the G@A site than at the GP300 site, a shower with the same number of particles will emit a weaker radio signal (i.e., a smaller-amplitude electric field) in G@A than at GP300. This leads to weaker signals in the antennas, making it generally more challenging to detect showers in G@A.
@@
Figure 4: GRAND detection unit deployed in GRAND@Auger. This is the result of converting an AERA station into a G@A detection unit. Left: The green pole is an aluminium tube attached to the original AERA pole with an adapter ring to increase the antenna height to 3 m (Section 3.1). The antenna head, containing the low-noise amplifier (Section 3.2), is connected to a sleeve on the top of the pole, which is kept in place with a bolt. A communications antenna is mounted midway along the pole. The triangular shaped box in the lower half, from the original AERA design, contains the battery, the front-end board (Section 3.3), and the charge controller. It also serves as support for the solar panel. Right: Photograph of a unit deployed at the G@A site, in Argentina. The plastic mesh around the DU protects it from free-roaming animals in the area.

### v2

@@ Abstract
The Giant Radio Array for Neutrino Detection (GRAND) is a proposed multi-messenger observatory of Ultra-High-Energy (UHE) particles of cosmic origin. Its main goal is to find the long-sought origin of UHE cosmic rays by detecting large numbers of them and the secondary particles created by their interactions like gamma rays and neutrinos. The GRAND Collaboration plans to achieve this using large arrays of radio antennas that look for the radio signals emitted by the air showers initiated by the interactions of the UHE particles in the atmosphere. Since 2023, three small-scale prototype GRAND arrays have been in operation: GRAND@Nançay in France, GRAND@Auger in Argentina, and GRANDProto300 in China. Together, their goal is to validate the detection principle of GRAND under prolonged field conditions, achieving efficient, autonomous radio-detection of air showers. We describe the hardware, software, layout, and operation of the GRAND prototypes. Using their data, we show a first characterization of the local electromagnetic environment of each site and a measurement of the Galactic synchrotron emission. Despite challenges, the successful operation of the prototypes confirms that the GRAND instrumentation is apt to address the goals of the experiment and lays the groundwork for its ensuing stages.
@@ 1 Introduction
The Giant Radio Array for Neutrino Detection (GRAND) [11] is a planned multi-messenger observatory of Ultra-High Energy (UHE) particles of cosmic origin—cosmic rays, neutrinos, and gamma rays. Its ultimate goal is to provide crucial evidence that addresses the long-standing question of the origin of UHE Cosmic Rays (UHECRs) [14, 12, 22]. GRAND plans to do so, directly, by detecting a vast number of UHECRs and, indirectly, by detecting secondary gamma rays and UHE neutrinos produced from the interaction of UHECRs [19]. To achieve this, GRAND plans to use large arrays of radio antennas that search for distinctive radio frequency signals emitted by extensive air showers initiated by UHE particles interacting in the atmosphere [28, 49].
@@
The radio-detection of UHECRs is a mature technique. It has been demonstrated in past and present detectors using overground antenna arrays [28, 49] by the Auger Engineering Radio Array (AERA) [7, 5], the Low-Frequency Array (LOFAR) [52, 48], the Taiwan Astroparticle Radiowave Observatory for Geo-synchrotron Emissions (TAROGE) [54], the Tianshan Radio Experiment for Neutrino Detection (TREND) [20], and the Owens Valley Radio Observatory Long Wavelength Array (OVRO-LWA) [38]; using in-ice arrays [18] by the Askaryan Radio Array (ARA) [10] and the Antarctic Ross Ice Shelf Antenna Neutrino Array (ARIANNA) [17]; and using balloon flights by the Antarctic Impulsive Transient Antenna (ANITA) [23, 26]. Concurrent with GRAND, the new Radio Detector of the Auger Prime upgrade [27], the upcoming Payload for Ultrahigh Energy Observations (PUEO) [4], and the Radio Neutrino Observatory in Greenland (RNO-G) [9], as well as the proposed next-generation Beamforming Elevated Array for COsmic Neutrinos (BEACON) [50], Radar Echo Telescope for Cosmic Rays (RET-CR) [45], and POEMMA-Balloon with Radio (PBR) [21] also target the radio-detection of UHE particles, though adopting different, complementary [55] detection strategies. References [34, 8, 24] present an overview of the upcoming UHE detectors.
@@
Figure 1: The GRAND prototype arrays. The three GRAND prototype arrays were deployed and have been in operation since 2023. They share a common basic instrumentation design, modified to suit the local environmental condition at the array site, and to test the performance of alternative hardware and software choices. This BlankMap-World-noborders.png has been obtained by the author(s) from the Wikimedia website, where it is stated to have been released into the public domain. It is included within this article on that basis (https://commons.wikimedia.org/wiki/File:BlankMap-World-noborders.png).
@@
Building on the experience gained by these experiments—most directly, by AERA and TREND—GRAND plans to deploy antenna arrays at a larger scale [11, 31]. This involves increasing the number of deployed antennas—to eventually tens of thousands—and deploying them in arrays that are sparser than before—with km-scale antenna separations. This up-scaling is required for the antenna arrays to monitor the large volumes of the atmosphere needed, creating an aperture of 107000 km2sr, which delives vast numbers of detected UHECRs—6500 UHECRs above $10^{19.5}$ eV per year [11]—and a sensitivity to even tiny fluxes of UHE neutrinos—2 orders of magnitude below the IceCube sensitivity [11, 37, 8, 34, 24].
@@
We implemented a staged construction strategy for GRAND [11, 30, 39] intended to validate the design and operation of the experiment on smaller scales before deploying larger arrays. Figure 1 shows the locations of the GRAND prototypes, in operation since 2023 [39]: GRAND@Nançay in Nançay, France, GRANDProto300 [44] in Dunhuang, China, and GRAND@Auger [16] in Malargüe, Argentina, at the site of the Pierre Auger Observatory (Pierre Auger) [3].
@@
Their purpose is to validate two main requirements that GRAND must fulfill to achieve its goals. First, the prototypes must validate the viability of autonomous radio-detection of air showers in a sparse radio array [40, 47, 42, 41], unaided by particle detectors, and despite the presence of stochastic noise and anthropogenic background that may mimic the electromagnetic signatures expected from showers. Second, the prototypes must validate that the arrays are able to operate under harsh environmental field conditions—high temperature, sandstorms, rainstorms—for prolonged periods of time. The lessons gained from the prototypes inform the design of the future stages of the experiment.
@@
In this paper, we introduce the design and operation of the GRAND prototype arrays, focusing on the two larger ones, GRANDProto300 (GP300) and GRAND@Auger (G@A), and we present a first look at their data. The remainder of the paper is organized as follows. Section 2 introduces the GRAND prototype arrays, their sites, and layouts. Section 3 describes the GRAND detection unit, its hardware, and performance. Section 4 presents the self-trigger algorithm and the software processes that comprise the GRAND data acquisition chain. Section 5 describes how the collected data are managed, i.e., the data transfer pipeline from local and central storage on-site to permanent storage off-site, and the data-quality monitoring tools of the experiment. Section 6 shows the first characterization of the local electromagnetic environment and the measurement of the Galactic synchrotron emission using data from the GRAND prototypes. Section 7 summarizes and presents the perspectives.
@@ 2 Overview of the GRAND prototype arrays
Figure 2: Array layout of GRANDProto300 (left) and of GRAND@Auger (right). Each GRAND Detection Unit (DU) has an assigned identification number. The Central Station in GRANDProto300 and the Central Radio Station in GRAND@Auger house the central data acquisition (DAQ) systems of the experiments. Positions for GRANDProto300 are given with respect to DU 1 of GP300_s65. For GRAND@Auger, DU positions are given in the Auger Coordinate System. We also show the nearby AERA radio-stations (in grey) and access roads.
The site and array layout for each GRAND prototype were decided to fulfill their specific goals: testbenching for GRAND@Nançay, demonstrating the autonomous radio-detection of air showers for GP300, and observing coincident detection of air showers with the Pierre Auger detectors in G@A. The radio-frequency electromagnetic environment varies between sites, with the anthropogenic background at GRAND@Nançay being the loudest. For GP300, the background is appreciably smaller than at G@A, predominantly due to GP300 being farther from human activity.
GRAND@Nançay is a small array of four GRAND Detection Units (DUs) located at the Nançay Radio Observatory, in France. It was conceived as a field site accessible to European laboratories, facilitating testbenching of equipment and new ideas in design, trigger, and data acquisition. The design of GRAND@Nançay resembles that of the other two prototype arrays, though with differences, such as having antennas of lower height and communications via optical fiber instead of wireless. Its small size is sufficient for its purpose as a testbed. In the remainder of the paper, we focus on GRANDProto300 and GRAND@Auger.
GRANDProto300 is the GRAND pathfinder array in the Northern Hemisphere. In its final stage, it will be a 300-antenna array; currently, it exists in partial configurations.
@@
In June 2023, the deployment of the 13-antenna stage, GP300_s13, was completed [43]. During its nearly two years of operation, it enabled the validation of the GRAND detection concept and the improvement of the system design. By April 2025, the array was extended to its current 65-antenna stage, GP300_s65, which has been operating since, and whose additional goals include testing the efficiency and purity of the autonomous radio-detection of inclined air showers from UHECRs. GP300, even in partial configuration, is expected to detect UHECRs with energies between 100 PeV and a few EeV, probing the Galactic to extragalactic transition region of cosmic ray origin.
@@
Figure 2 shows the layout of GP300_s13 and GP300_s65. In both GP300 and G@A, DUs are arranged in a hexagonal layout, a choice made to optimize the simulated efficiency of air-shower detection with respect to the density of antennas on the ground. A hexagonal array has fewer directions along which the performance of air shower-reconstruction methods is degraded as a result of the shower trajectory being aligned with the axes of symmetry of the array layout.
@@
GP300_s13 had a three-hexagon layout, covering an area of about 2 km2 with a DU spacing of about 500 m. The Central Station 1 (CS1) was built north-west of the array; it consisted of three sets of container rooms that stored the central data acquisition (DAQ) computer and the lodging and cooking facilities. GP300_s13 was decommissioned in April 2025 and its DUs were incorporated into GP300_s65, which has remained operational since. GP300_s65 was built south of GP13, covering an area of about 30 km2. The array has a 55-DU dense region to the east, with a DU spacing of 577 m, and a 10-DU sparse region to the west, with a spacing of 1000 m. A new Central Station (CS2) was built for GP300_s65 and placed in its center.
@@
Preliminary simulations [46] estimate that the full GP300 array may detect about 130 inclined cosmic-ray events per day—for an ideal 100% uptime of the detector—in the energy range of $10^{17}$$10^{20}$ eV, reaching a trigger efficiency of 90% already at $10^{18.3}$ eV. These estimates account for the processing of the radio signal from the air shower through the Radio-Frequency (RF) chain of the detector units and for the persistent noise from the Galactic emission. Improvements are underway to optimize the cosmic-ray trigger rate, suppress the electromagnetic noise and background, and refine the selection of cosmic-ray candidates [29, 33, 2, 25, 1, 32]. Because GP300_s13 and GP300_s65 are smaller, they are sensitive to the lower end of the cosmic-ray energy range, between about $10^{16.5}$ eV and $10^{18.5}$ eV, where the cosmic-ray flux is higher. The search for these cosmic rays has already yielded its first likely candidates in GP300_s65 data [29, 33, 2].
@@
GRAND@Auger results from an agreement between the GRAND and Pierre Auger Collaborations to repurpose 10 AERA Phase-2 radio-stations in Malargüe, Argentina, into GRAND DUs. The G@A array is embedded within the region of the Pierre Auger detector array, consisting of water-Cherenkov particle detectors (Surface Detectors—SD), fluorescence telescopes, and radio antennas [27]. This offers, uniquely among the three GRAND prototypes, the opportunity to validate the radio-detection—and the ensuing reconstruction—of air showers that are simultaneously detected by both the Pierre Auger Observatory and the GRAND@Auger antennas.
@@
The deployment of G@A was completed in August 2023 and it has been operational since then. Figure 2, right, shows the layout of G@A. The GRAND prototype was deployed within the AERA array, itself located within the Pierre Auger 750-m SD array, and within the field of view of the Coihueco and HEAT fluorescence detectors [6, 36]. The G@A DUs are laid on two superimposed hexagons with a predominant spacing of 250 m, covering an area of 0.5 km2. The central DAQ computer is located in the AERA Central Radio Station (CRS). Nearby roads facilitate access to the site.
@@
For G@A, using a denser detector layout enhances the detection of less energetic and less inclined events. The estimated event rate for the Pierre Auger 750-m SD array is $390\pm 17$ showers per day over 25 km2 [35]. This corresponds to about 8 showers per day detected by the Pierre Auger SD array within the area of G@A, making it viable to search for showers measured simultaneously by G@A and Pierre Auger detector arrays. One such coincident event has been detected so far [16]. Because the geomagnetic field is weaker at the G@A site than at the GP300 site, a shower with the same number of particles will emit a weaker radio-frequency signal (i.e., a smaller-amplitude electric field) in G@A than at GP300. This leads to weaker signals in the antennas, making it generally more challenging to detect showers in G@A.
@@
Figure 4: GRAND detection unit deployed in GRAND@Auger. Left: The green pole is an aluminum tube attached to the original AERA pole with an adapter ring to increase the antenna height to 3 m (Section 3.1). The antenna head, containing the low-noise amplifier (Section 3.2), is connected to a sleeve on the top of the pole, which is kept in place with a bolt. A communications antenna is mounted midway along the pole. The triangular shaped box in the lower half, from the original AERA design, contains the battery, the front-end board (Section 3.3), and the charge controller. It also serves as support for the solar panel. Right: Photograph of a unit deployed at the G@A site, in Argentina. The plastic mesh around the DU protects it from free-roaming animals in the area.